OpenStudy (anonymous):

evaluate the function at the given values of x, if possible f(x)=6x^2-1/x^2 (a)x=2 (b)x=-1

OpenStudy (anonymous):

a) 24-1/4=95/4 b)5

OpenStudy (anonymous):

\[f(2) = 6(2^2) - {1 \over 2^2} = 6*4 - {1 \over 4} = 24 - 0.25 = 23.75\] \[f(-1) = 6(-1)^2 - {1 \over (-1)^2} = 6 - {1 \over 1} = 6 - 1 = 5\]